Suppose, in an economy, currency in circulation (C) is $16 billions, reserves (R) held by banks are $4
billions, and deposits (D) by people and firms in banks are worth $ 84 billions. If there are no excess
reserves, then
(a) What is the money supply (M) in the economy? _______________

(b) What is the monetary base (MB)? _______________

(c) What is the currency deposit ratio ? _______________

(d) What is the reserve deposit ratio? _______________

(e) What is the money multiplier (m)? _______________

(f) If the monetary base increases by 10%, then the money supply increases by $ _________
billions.

(g) In general, how can the central bank influence the monetary base?
Answer #1

Answer

(a) M = C + D = 16 billion + 84 billion= $100 billion

(b) MB = C + R = 16 billion + 4 billion = $20 billion

(c) currency deposit ratio(cr) = C/D = (16 billion) / (84 billion) = 4/21 = 0.19(approx)

(d) Reserve deposit rate(rr) = RR/D = (4 billion)/(84 billion) = 1/21 = 0.0476(approx) Here there is no excess reserves and hence Required reserves(RR) = Total reserves(R)

(e) Money multiplier(m) = (1 + cr)/(cr + rr + er) where er = excess reserve ratio = 0.

=> Money multiplier = (1 + 4/21)/(4/21 + 1/21) = 5

(f) 10% of monetary base = 10% of 20 billion = 2 billion.

Thus new monetary base = 22 billion.

Money multiplier(m) is 5.

Thus New Money supply = m*MB = 22 billion*5 = 110 billion.

Thus money supply will increase by 110 billion - 100 billion = $10 billion

(g) Central bank influences Monetary base by changing reserves for the banks and they do this by using open market sale and purchase of bonds.